A gas occupies 40 mL at -123C. What Volume does it occupy at 27C, assuming pressure is constant?
1
Expert's answer
2020-04-27T01:53:33-0400

vRT=PV law of ideal gas

because, v- nuumber of moles, R - constant, Pressure - constant

T_1/V_1=T_2/V_2, temperature in Kelvin.

T_1=-123°C=150K, Volume_1=40 ml

T_2=27°C=300K, Volume_2=?

V_2=T_2 * V_1/T_1=40ml * 300K/150K=80ml